# Mechanistic Indicators of Understanding in Large Language Models

### v4

@@ 5  A Motley Mix of Mechanisms
Yet humans possess a distinctive meta-cognitive ability to override heuristics and unify phenomena using increasingly parsimonious explanatory structures. This striving for parsimony is widely regarded as a hallmark of world understanding. As Einstein put it, the “grand aim of all science … is to cover the greatest possible number of empirical facts by logical deductions from the smallest possible number of hypotheses or axioms” (quoted in Nash 1963, p. 173). It is this mature understanding—marked by a striving for parsimony, where a few principles unify a wide range of phenomena—that supports robust explanation and prediction and ultimately warrants epistemic trust (de Regt, 2009; Queloz and Beckmann, 2026).
The motley mix phenomenon, however, undermines this basis for epistemic trust: even broad behavioral success cannot reveal whether the model relies on parsimonious unifying mechanisms or merely a swarm of shallow heuristics. Moreover, even when researchers do identify circuits implementing principled, parsimonious understanding, these understanding-bearing mechanisms might still be overwhelmed by other heuristics. Researchers observed this problem in syllogistic inference: they identified a formal circuit that correctly processes logical structure, but gets drowned out by content-based mechanisms (Kim et al., 2025; Valentino et al., 2025). For instance, the model is less likely to accept the syllogism: “All apples are vegetations. All vegetations are institutions. Some apples are institutions” than “All apples are edible fruits. All edible fruits are fruits. All apples are fruits”—despite both being logically valid—simply because the first contains substantively implausible premises that triggers content-based circuits. Because whatever genuine understanding an LLM possesses is buried within a sprawl of shallower heuristics, it cannot ground the same epistemic trust we grant to more parsimonious minds.
@@ 6  Conclusion
Yet this investigation has also revealed a profound alienness. LLMs do not strive for parsimony as humans do; whatever understanding they possess is distributed across a motley mix of mechanisms. Recognizing this need not undermine the claim that LLMs understand, but it does require us to draw finer distinctions between varieties of understanding and to forge conceptions that are decoupled from the specificities of human understanding. [^34] For discussions of this need for new ways of thinking and talking about LLMs, see Shanahan 2024b; 2024a and Müller & Löhr 2026.

### v5

@@ 5  A Motley Mix of Mechanisms
Yet humans possess a distinctive meta-cognitive ability to override heuristics and unify phenomena using increasingly parsimonious explanatory structures. This striving for parsimonious systematicity is widely regarded as a hallmark of world understanding (Queloz, 2025a, ; Queloz, 2025b, ). As Einstein put it, the “grand aim of all science … is to cover the greatest possible number of empirical facts by logical deductions from the smallest possible number of hypotheses or axioms” (quoted in Nash 1963, p. 173). It is this mature understanding—marked by a striving for parsimony, where a few principles unify a wide range of phenomena—that supports robust explanation and prediction and ultimately warrants epistemic trust (de Regt,, 2009; Queloz and Beckmann,, 2026).
The motley mix phenomenon, however, undermines this basis for epistemic trust: even broad behavioral success cannot reveal whether the model relies on parsimonious unifying mechanisms or merely a swarm of shallow heuristics. Moreover, even when researchers do identify circuits implementing principled, parsimonious understanding, these understanding-bearing mechanisms might still be overwhelmed by other heuristics. Researchers observed this problem in syllogistic inference: they identified a formal circuit that correctly processes logical structure, but gets drowned out by content-based mechanisms (Kim et al.,, 2025; Valentino et al.,, 2025). For instance, the model is less likely to accept the syllogism: “All apples are vegetations. All vegetations are institutions. Some apples are institutions” than “All apples are edible fruits. All edible fruits are fruits. All apples are fruits”—despite both being logically valid—simply because the first contains substantively implausible premises that triggers content-based circuits. Because whatever genuine understanding an LLM possesses is buried within a sprawl of shallower heuristics, it cannot ground the same epistemic trust we grant to more parsimonious minds.
@@ 6  Conclusion
Yet this investigation has also revealed a profound alienness. LLMs do not strive for parsimony as humans do; whatever understanding they possess is distributed across a motley mix of mechanisms. Recognizing this need not undermine the claim that LLMs understand, but it does require us to draw finer distinctions between varieties of understanding and to forge conceptions that are decoupled from the specificities of human understanding. [^34] For discussions of this need for new ways of thinking and talking about LLMs, see Shanahan ; and Müller & Löhr 2026.